# Fast relaxation of a viscous vortex in an external flow

### v1

@@ 3.1 The short time estimate
Altogether, since $\varepsilon^{2}/\delta=t/T_{0}$, we have shown that
$\frac{1}{\delta}\,\bigl{|}\langle w,\mathcal{A}[w]\rangle_{\mathcal{Y}}\bigr{|}\,\leq\,\frac{t}{T_{0}}\,\bigl{(}\mathcal{F}[w]+C\mathcal{E}[w]\bigr{)}\,.$ (3.13)
@@ 3.2 Construction of the energy functional
The approach of the previous section is relatively simple and provides a control of the solution of (3.2) in the natural function space $\mathcal{Y}$. However, as can be seen from the left-hand side of (3.14), the argument completely breaks down when $t/T_{0}>\kappa$. To reach longer times, it is necessary to use a more sophisticated energy functional which allows us to treat separately three regions of the physical space: a small neighborhood of the vortex center, an intermediate region, and a far field region where the influence of the vortex is negligible. This idea is implemented in the previous work [4], which deals with the interaction of localized vortices. In this section, we provide a simplified version of the argument, which gives slightly stronger results.
@@
$q_{\varepsilon}(\xi,t)\,=\,\frac{|\xi|^{2}}{4}+\frac{\varepsilon^{2}T_{0}}{4v_{0}(\xi)}\,\Bigl{(}b(t)(\xi_{1}^{2}-\xi_{2}^{2})-2a(t)\xi_{1}\xi_{2}\Bigr{)}\,,\qquad\forall\,(\xi,t)\in\mathbb{R}^{2}\times(0,T)\,,$ (3.18)
@@
$v_{0}(\xi)\,=\,\frac{1}{2\pi|\xi|^{2}}\Bigl{(}1-e^{-|\xi|^{2}/4}\Bigr{)}\,,\qquad\forall\,\xi\in\mathbb{R}^{2}\,.$ (3.19)
@@
$\rho(\theta,t)\,=\,\pi T_{0}\bigl{(}a(t)\sin(2\theta)-b(t)\cos(2\theta)\bigr{)}A^{2}+\mathcal{O}(A^{4})\,,\qquad\text{as }\,A\to 0\,.$ (3.23)
@@
provided $A$ and $\varepsilon$ are small enough. This implies that the function $g_{\varepsilon}$ is strictly increasing on the interval $[0,A^{2}/\varepsilon^{2}]$ with $g_{\varepsilon}(0)=0$ and $g_{\varepsilon}\bigl{(}A^{2}/\varepsilon^{2})\geq A^{2}/(2\varepsilon^{2})$. By the intermediate value theorem, the equation $g_{\varepsilon}(s)=A^{2}/(4\varepsilon^{2})$ has a unique solution $s=\bar{s}(\theta,t)$ in that interval, and the implicit function theorem ensures that $\bar{s}(\theta,t)$ is a smooth function of $\theta$ and $t$. Moreover, we easily deduce from (3.25) that $A^{2}/(6\varepsilon^{2})\leq\bar{s}\leq A^{2}/(2\varepsilon^{2})$. If we assume that $\varepsilon=\mathcal{O}(A^{\alpha})$ for some $\alpha>1$, this implies that $\bar{s}=\mathcal{O}(A^{2-2\alpha})$, hence $\phi(\bar{s})=\bar{s}^{2}+\mathcal{O}(A^{\infty})$. Returning to (3.24), we deduce that
@@
According to the definition (3.21), the weight function $p_{\varepsilon}(\xi,t)$ is equal to $\exp\bigl{(}q_{\varepsilon}(\xi,t)\bigr{)}$ in the (elliptical) inner region $\mathrm{I}_{\varepsilon}(t)$, at the boundary of which it takes the value $\exp\bigl{(}A^{2}/(4\varepsilon^{2})\bigr{)}$ by construction. It is then extended as a constant function in the intermediate (annular) region $\mathrm{II}_{\varepsilon}(t)$, and as a Gaussian function in the exterior region $\mathrm{III}_{\varepsilon}(t)$. Since $\gamma=A^{2}/B^{2}$, we observe that $\exp(\gamma|\xi|^{2}/4)=\exp\bigl{(}A^{2}/(4\varepsilon^{2})\bigr{)}$ when $|\xi|=B/\varepsilon$, which implies that $p_{\varepsilon}(\xi,t)$ is a Lipschitz continuous function of $\xi\in\mathbb{R}^{2}$. We also have uniform bounds of the form
@@ 3.3 The large time estimate
The goal of this section is to prove Proposition 3.3. If $w(\cdot,t)$ is the solution of (3.2) with zero initial data, a direct calculation shows that the energy function (3.29) satisfies, as in (3.9),
@@
In (3.34) it is understood that $\varepsilon=\sqrt{\nu t}/d$ as usual, so that $t\partial_{t}\varepsilon=\varepsilon/2$. Except for that relation, we can consider the quantities introduced in (3.35) as defined for any fixed $t\in(0,T)$ and for an arbitrary value of the (small) parameter $\varepsilon$. Useful estimates on these quantities are derived in the following paragraphs.
@@ 3.3.1 Control of the diffusion terms
The quadratic form $\mathcal{Q}_{\varepsilon,t}$ is everywhere coercive except in the region $\mathrm{II}_{\varepsilon}$ where the weight function $p_{\varepsilon}$ is independent of $\xi$. The following lower bound can be established as in [6, Prop. 4.15]. For the reader’s convenience, we provide the details in Section A.4.
@@
In view of (3.36) and (3.38), what remains is to estimate the time derivative of the weight function $p_{\varepsilon}$. In the region $\mathrm{I}_{\varepsilon}$ we have $p_{\varepsilon}=\exp(q_{\varepsilon})$ where $q_{\varepsilon}$ is given by (3.18). Recalling that $t\partial_{t}\varepsilon^{2}=\varepsilon^{2}$, we find
$t\partial_{t}q_{\varepsilon}(\xi,t)\,=\,\frac{\varepsilon^{2}T_{0}}{4v_{0}(\xi)}\,\Bigl{(}\bigl{[}b(t)+tb^{\prime}(t)\bigr{]}(\xi_{1}^{2}-\xi_{2}^{2})-2\bigl{[}a(t)+ta^{\prime}(t)\bigr{]}\xi_{1}\xi_{2}\Bigr{)}\,.$ (3.40)
@@
where the constant $C$ only depends on the quantity $\mathcal{K}$ defined in (1.9). Since $|v_{0}(\xi)|\leq C/(1+|\xi|^{2})$ and $\varepsilon^{2}|\xi|^{2}\leq 2A^{2}$ in region $\mathrm{I}_{\varepsilon}$, we deduce from (3.40), (3.41) that
@@
As is easily verified, we have $|\xi|^{2}\leq 4\chi_{\varepsilon}$ in region $\mathrm{I}_{\varepsilon}$, so that the quantity inside square brackets is larger than $\kappa(\chi_{\varepsilon}+1)/2$ if $A>0$ is small enough. Similarly $\chi_{\varepsilon}\leq A^{2}/\varepsilon^{2}$ in region $\mathrm{II}_{\varepsilon}$, which implies that $A^{2}/(8\varepsilon^{2})-1\geq(\chi_{\varepsilon}+1)/16$ if $\varepsilon\ll A$. So, assuming that $\kappa\leq 1/16$, we find
@@ 3.3.2 Control of the advection terms
where $U_{0}(\xi)=\xi^{\perp}v_{0}(\xi)$, $E_{2}$ is given by (2.9), $|\hat{U}_{2}(\xi,t)|\leq C/(1+|\xi|)$, and $|\bar{E}_{3}(\xi,t)|\leq C(1+|\xi|^{2})$. To estimate the quantity $\mathcal{A}_{\varepsilon,t}^{(1)}[w]$, we first assume that $\xi\in\mathrm{I}_{\varepsilon}$, so that $p_{\varepsilon}=\exp(q_{\varepsilon})$. Using the explicit expression (3.18), we find by a direct calculation
$U_{0}\cdot\nabla q_{\varepsilon}\,=\,\frac{\varepsilon^{2}T_{0}}{4}\,\xi^{\perp}\cdot\nabla\Bigl{(}b(\xi_{1}^{2}-\xi_{2}^{2})-2a\xi_{1}\xi_{2}\Bigr{)}\,=\,-\frac{\varepsilon^{2}T_{0}}{2}\Bigl{(}a(\xi_{1}^{2}-\xi_{2}^{2})+2b\xi_{1}\xi_{2}\Bigr{)}\,=\,-\frac{\varepsilon^{2}}{2}\,\xi\cdot E_{2}\,,$
where the last equality follows from (2.9). We thus have a partial cancellation between the terms $U_{0}\cdot\nabla q_{\varepsilon}$ and $\varepsilon^{2}E_{2}\cdot\nabla q_{0}$, which is the reason for which we included a non-radially symmetric correction in the definition (3.18) of the function $q_{\varepsilon}$. It follows that
$\bigl{(}U_{\mathrm{app}}+E(f,z)\bigr{)}\cdot\nabla q_{\varepsilon}\,=\,\varepsilon^{2}E_{2}\cdot\nabla\Bigl{(}q_{\varepsilon}-\frac{|\xi|^{2}}{4}\Bigr{)}+\varepsilon^{2}\hat{U}_{2}\cdot\nabla q_{\varepsilon}+\varepsilon^{3}\bar{E}_{3}\cdot\nabla q_{\varepsilon}\,.$
Since $|\nabla q_{\varepsilon}|\leq C|\xi|$ and $|\nabla(q_{\varepsilon}-|\xi|^{2}/4)|\leq C\varepsilon^{2}|\xi|(1+|\xi|^{2})$ for $\xi\in\mathrm{I}_{\varepsilon}$, we easily obtain
@@
$\int_{\{|\xi|\leq\varepsilon^{-1/2}\}}\bigl{|}p_{\varepsilon}-p_{0}\bigr{|}\,|w|\,|v|\,|\nabla\Omega_{0}|\,\mathrm{d}\xi\,\leq\,C\varepsilon^{2}\,\int_{\mathrm{I}_{\varepsilon}}|\xi|^{3}(1+|\xi|^{2})\,|w|\,|v|\,\mathrm{d}\xi\,\leq\,C\varepsilon^{2}\,\mathcal{E}_{\varepsilon,t}[w]\,.$
@@
In the previous paragraphs we obtained accurate bounds on the diffusion terms $\mathcal{D}_{\varepsilon,t}[w]$ and the advection term $\mathcal{A}_{\varepsilon,t}[w]$ in (3.35). As for the nonlinear term $\mathcal{N}_{\varepsilon,t}[w]$ and the source term $\mathcal{S}_{\varepsilon,t}[w]$, they can be estimated exactly as in the proof of Proposition 3.1. We thus find
@@
On the other hand, we have $\Omega_{\mathrm{app}}-\Omega_{0}-\varepsilon^{2}\bar{\Omega}_{2}=\delta\varepsilon^{2}\tilde{\Omega}_{2}+\varepsilon^{3}\Omega_{3}+\varepsilon^{4}\Omega_{4}$, as can be seen from the expansion (2.5) and the definition of $\Omega_{2}$ in Section 2.3. We deduce that $\mathcal{I}_{2}(t)\leq C\varepsilon^{2}(\varepsilon+\delta)$, and together with (3.48) this concludes the proof of estimate (1.13). ∎
@@ 3.3.4 Alternative definitions of the vortex position
The right-hand side can be estimated using Proposition 3.3, which gives
@@
On the other hand, the approximate vortex solution $\hat{z}(t)$ given by the simple ODE (1.8) is only $\mathcal{O}(\varepsilon^{2})$ close to the solution $z(t)$ of (1.14), unless of course the additional term $\Delta f(z(t),t)$ in (1.14) vanishes identically. As was already mentioned, this is the case if the external velocity field $f$ is irrotational, see the discussion in Remark 1.1. In general, taking the difference of (1.8) and (1.14), we obtain the inequality
@@ 3.3.5 Alternative choice of the weight function
The estimates are even simpler because the intermediate region, where the dangerous advection terms give no contribution, is now larger. Although somewhat weaker, the result is still sufficient to imply Theorem 1.6.
@@ 4.1 Enhanced dissipation estimates
$|g_{0}(\xi)|\,\leq\,\chi(\xi)\,e^{|\xi|^{2}/8}|\phi_{0}(\xi)|\,\leq\,\chi(\xi)\,e^{-(\gamma-1/8)|\xi|^{2}}\|\phi_{0}\|_{\gamma}\,\leq\,e^{-\mu|\xi|^{2}}\Bigl{(}\chi(\xi)\,e^{-\mu|\xi|^{2}}\Bigr{)}\|\phi_{0}\|_{\gamma}\,,$
@@
Applying (4.10) and Hölder’s inequality, we thus obtain

### v2

@@ 3.1 The short time estimate
Altogether, recalling that $\varepsilon^{2}/\delta=t/T_{0}$, we find
$\frac{1}{\delta}\,\bigl{|}\langle w,\mathcal{A}[w]\rangle_{\mathcal{Y}}\bigr{|}\,\leq\,\frac{t}{T_{0}}\,\bigl{(}\mathcal{F}[w]+C\mathcal{E}[w]\bigr{)}\,,\qquad\text{for some }C>0\,.$ (3.13)
@@ 3.2 Construction of the energy functional
The approach of the previous section is relatively simple and provides a control of the solution of (3.2) in the natural function space $\mathcal{Y}$. However, as can be seen from the left-hand side of (3.14), the argument completely breaks down when $t/T_{0}>\kappa$. To reach longer times, it is necessary to use a more sophisticated energy functional which allows us to treat separately three regions of the physical space: a small neighborhood of the vortex center, an intermediate region, and a far field region where the influence of the vortex is negligible. This idea is implemented in the previous work [4], where the interaction of localized vortices is studied. In this section, we provide a simplified version of the argument, which gives slightly stronger results.
@@
$q_{\varepsilon}(\xi,t)\,=\,\frac{|\xi|^{2}}{4}+\frac{\varepsilon^{2}T_{0}}{4v_{*}(|\xi|)}\,\Bigl{(}b(t)(\xi_{1}^{2}-\xi_{2}^{2})-2a(t)\xi_{1}\xi_{2}\Bigr{)}\,,\qquad\forall\,(\xi,t)\in\mathbb{R}^{2}\times(0,T)\,,$ (3.18)
@@
$v_{*}(|\xi|)\,=\,\frac{1}{2\pi|\xi|^{2}}\Bigl{(}1-e^{-|\xi|^{2}/4}\Bigr{)}\,,\qquad\forall\,\xi\in\mathbb{R}^{2}\,.$ (3.19)
@@
$\rho(\theta,t)\,=\,\pi T_{0}\bigl{(}a(t)\sin(2\theta)-b(t)\cos(2\theta)\bigr{)}A^{2}+\mathcal{O}(A^{4})\,.$ (3.23)
@@
provided $A$ and $\varepsilon$ are small enough. This implies that the function $g_{\varepsilon}$ is strictly increasing on the interval $[0,A^{2}/\varepsilon^{2}]$ with $g_{\varepsilon}(0)=0$ and $g_{\varepsilon}\bigl{(}A^{2}/\varepsilon^{2})\geq A^{2}/(2\varepsilon^{2})$. By the intermediate value theorem, the equation $g_{\varepsilon}(s)=A^{2}/(4\varepsilon^{2})$ has a unique solution $s=\bar{s}(\theta,t)$ in that interval, and the implicit function theorem ensures that $\bar{s}(\theta,t)$ is a smooth function of $\theta$ and $t$. Moreover, we easily deduce from (3.25) that $A^{2}/(6\varepsilon^{2})\leq\bar{s}\leq A^{2}/(2\varepsilon^{2})$. If we assume that $\varepsilon=\mathcal{O}(A^{\alpha})$ for some $\alpha>1$, this implies that $\phi(\bar{s})=\bar{s}^{2}+\mathcal{O}(A^{\infty})$. Returning to (3.24), we deduce that
@@
According to the definition (3.21), the weight function $p_{\varepsilon}(\xi,t)$ is equal to $\exp\bigl{(}q_{\varepsilon}(\xi,t)\bigr{)}$ in the (elliptical) inner region $\mathrm{I}_{\varepsilon}(t)$, at the boundary of which $p_{\varepsilon}(\xi,t)=\exp\bigl{(}A^{2}/(4\varepsilon^{2})\bigr{)}$ by construction. The weight $p_{\varepsilon}$ is then extended as a constant function in the intermediate (annular) region $\mathrm{II}_{\varepsilon}(t)$, and as a Gaussian function in the exterior region $\mathrm{III}_{\varepsilon}(t)$. Since $\gamma=A^{2}/B^{2}$, we observe that $\exp(\gamma|\xi|^{2}/4)=\exp\bigl{(}A^{2}/(4\varepsilon^{2})\bigr{)}$ when $|\xi|=B/\varepsilon$, which means that $p_{\varepsilon}$ has no jump at the common boundary of $\mathrm{II}_{\varepsilon}(t)$ and $\mathrm{III}_{\varepsilon}(t)$, and is therefore a Lipschitz continuous function of $\xi\in\mathbb{R}^{2}$. It is not difficult to verify that $p_{\varepsilon}$ satisfies uniform bounds of the form
@@ 3.3 The large time estimate
The goal of this section is to prove Proposition 3.3. If $w(\cdot,t)$ is the solution of (3.2) with zero initial data, a direct calculation shows that the energy function (3.29) satisfies
@@
In (3.34) it is understood that $\varepsilon=\sqrt{\nu t}/d$ as usual, so that $t\partial_{t}\varepsilon=\varepsilon/2$. Except for that relation, we can consider the quantities introduced in (3.35) as defined for any fixed $t\in(0,T)$ and any fixed $\varepsilon$ such that $0<\varepsilon\ll 1$. Useful estimates on these quantities are derived in the following paragraphs.
@@ 3.3.1 Control of the diffusion terms
The quadratic form $\mathcal{Q}_{\varepsilon,t}$ is everywhere coercive except in the region $\mathrm{II}_{\varepsilon}$ where $\nabla p_{\varepsilon}=0$. The following lower bound can be established as in [6, Prop. 4.15]. For the reader’s convenience, we provide the details in Section A.4.
@@
In view of (3.36) and (3.38), it remains to estimate the time derivative of the weight function $p_{\varepsilon}$. In the region $\mathrm{I}_{\varepsilon}$ we have $p_{\varepsilon}=\exp(q_{\varepsilon})$ where $q_{\varepsilon}$ is given by (3.18). Recalling that $t\partial_{t}\varepsilon^{2}=\varepsilon^{2}$, we find
$t\partial_{t}q_{\varepsilon}(\xi,t)\,=\,\frac{\varepsilon^{2}T_{0}}{4v_{*}(|\xi|)}\,\Bigl{(}\bigl{[}b(t)+tb^{\prime}(t)\bigr{]}(\xi_{1}^{2}-\xi_{2}^{2})-2\bigl{[}a(t)+ta^{\prime}(t)\bigr{]}\xi_{1}\xi_{2}\Bigr{)}\,.$ (3.40)
@@
where the constant $C$ only depends on the quantity $\mathcal{K}$ defined in (1.9). Since $\varepsilon^{2}|\xi|^{2}\leq 2A^{2}$ in region $\mathrm{I}_{\varepsilon}$ and $|v_{*}(|\xi|)|\leq C/(1+|\xi|^{2})$ by (3.19), we deduce from (3.40), (3.41) that
@@
As is easily verified, we have $|\xi|^{2}\leq 4\chi_{\varepsilon}$ in region $\mathrm{I}_{\varepsilon}$, so that the quantity inside the square brackets is larger than $\kappa(\chi_{\varepsilon}+1)/2$ if $A>0$ is small enough. Similarly $\chi_{\varepsilon}\leq A^{2}/\varepsilon^{2}$ in region $\mathrm{II}_{\varepsilon}$, which implies that $A^{2}/(8\varepsilon^{2})-1\geq(\chi_{\varepsilon}+1)/16$ if $\varepsilon\ll A$. So, assuming that $\kappa\leq 1/16$, we find
@@ 3.3.2 Control of the advection terms
where $U_{0}(\xi)=\xi^{\perp}v_{*}(|\xi|)$, $E_{2}$ is given by (2.9), $|\hat{U}_{2}(\xi,t)|\leq C/(1+|\xi|)$, and $|\bar{E}_{3}(\xi,t)|\leq C(1+|\xi|^{2})$. To estimate the quantity $\mathcal{A}_{\varepsilon,t}^{(1)}[w]$, we first assume that $\xi\in\mathrm{I}_{\varepsilon}$, so that $p_{\varepsilon}=\exp(q_{\varepsilon})$. Using the explicit expression (3.18) and denoting $q_{0}(\xi)=|\xi|^{2}/4$, we find by a direct calculation
$U_{0}\cdot\nabla q_{\varepsilon}\,=\,\frac{\varepsilon^{2}T_{0}}{4}\,\xi^{\perp}\cdot\nabla\Bigl{(}b(\xi_{1}^{2}-\xi_{2}^{2})-2a\xi_{1}\xi_{2}\Bigr{)}\,=\,-\frac{\varepsilon^{2}T_{0}}{2}\Bigl{(}a(\xi_{1}^{2}-\xi_{2}^{2})+2b\xi_{1}\xi_{2}\Bigr{)}\,=\,-\varepsilon^{2}E_{2}\cdot\nabla q_{0}\,,$
where the last equality follows from (2.9). We thus have a partial cancellation between the terms $U_{\mathrm{app}}\cdot\nabla q_{\varepsilon}$ and $E(f,z)\cdot\nabla q_{\varepsilon}$, which is precisely the reason for which we included a non-radially symmetric correction in the definition (3.18) of the function $q_{\varepsilon}$. It follows that
$\bigl{(}U_{\mathrm{app}}+E(f,z)\bigr{)}\cdot\nabla q_{\varepsilon}\,=\,\varepsilon^{2}E_{2}\cdot\nabla\bigl{(}q_{\varepsilon}-q_{0}\bigr{)}+\varepsilon^{2}\hat{U}_{2}\cdot\nabla q_{\varepsilon}+\varepsilon^{3}\bar{E}_{3}\cdot\nabla q_{\varepsilon}\,.$
Since $|\nabla q_{\varepsilon}|\leq C|\xi|$ and $|\nabla(q_{\varepsilon}-q_{0})|\leq C\varepsilon^{2}|\xi|(1+|\xi|^{2})$ for $\xi\in\mathrm{I}_{\varepsilon}$, we easily obtain
@@
$\int_{\{|\xi|\leq\varepsilon^{-1/2}\}}\bigl{|}p_{\varepsilon}-p_{0}\bigr{|}\,|w|\,|v|\,|\nabla\Omega_{0}|\,\mathrm{d}\xi\,\leq\,C\varepsilon^{2}\,\int_{\mathrm{I}_{\varepsilon}}|\xi|^{3}(1+|\xi|^{2})\,|w|\,|v|\,\mathrm{d}\xi\,\leq\,C\varepsilon^{2}\,\mathcal{E}_{\varepsilon,t}[w]\,,$
@@
In the previous paragraphs we obtained accurate bounds on the diffusion terms $\mathcal{D}_{\varepsilon,t}[w]$ and the advection term $\mathcal{A}_{\varepsilon,t}[w]$ in (3.35). As for the nonlinear term $\mathcal{N}_{\varepsilon,t}[w]$ and the source term $\mathcal{S}_{\varepsilon,t}[w]$, they can be estimated exactly as in the proof of Proposition 3.1. Using in particular the uniform bounds (3.28) and the estimate $|\nabla p_{\varepsilon}|\leq C\chi_{\varepsilon}^{1/2}p_{\varepsilon}$, which is established as in Lemma 3.6, we find
@@
On the other hand, we have $\Omega_{\mathrm{app}}-\Omega_{0}-\varepsilon^{2}\bar{\Omega}_{2}=\delta\varepsilon^{2}\tilde{\Omega}_{2}+\varepsilon^{3}\Omega_{3}+\varepsilon^{4}\Omega_{4}$, as can be seen from the expansion (2.5) and the definition of $\Omega_{2}$ in Section 2.3.1. We deduce that $\mathcal{I}_{2}(t)\leq C\varepsilon^{2}(\varepsilon+\delta)$, and together with (3.48) this concludes the proof of estimate (1.13). ∎
@@ 3.3.4 Alternative definitions of the vortex position
The integral in the right-hand side can be estimated using Proposition 3.3, which gives
@@
On the other hand, as computed below, the approximate vortex position $\hat{z}(t)$ given by the simple ODE (1.8) is only $\mathcal{O}(\varepsilon^{2})$ close to the solution $z(t)$ of (1.14), unless the additional term $\Delta f(z(t),t)$ in (1.14) vanishes identically. As was already mentioned, this is the case if the external velocity field $f$ is irrotational, see the discussion in Remark 1.1. In general, taking the difference of (1.8) and (1.14), we obtain the inequality
@@ 3.3.5 Alternative choice of the weight function
The estimates are even simpler because the intermediate region, where the dangerous advection terms give no contribution, is now larger. Although somewhat weaker, the result obtained in this way is still sufficient to imply Theorem 1.6.
@@ 4.1 Enhanced dissipation estimates
$|g_{0}(\xi)|\,=\,\chi(\xi)\,e^{|\xi|^{2}/8}|\phi_{0}(\xi)|\,\leq\,\chi(\xi)\,e^{-(\gamma-1/8)|\xi|^{2}}\|\phi_{0}\|_{\gamma}\,\leq\,e^{-\mu|\xi|^{2}}\Bigl{(}\chi(\xi)\,e^{-\mu|\xi|^{2}}\Bigr{)}\|\phi_{0}\|_{\gamma}\,,$
@@
Using (4.10) and applying Hölder’s inequality, we thus obtain